OpenStudy (anonymous):

Help...Write an equation of the line containing the given point and perpendicular to the given line (0,8); 5x+7y=9

OpenStudy (anonymous):

First, take your given line 5x+7y=9 and put it into slope - intercept form (y=mx+b) Second, determine the slope and use the OPPOSITE RECIPROCAL Then, plug your point and your NEW slope into point slope form y-y1=m(x-x1) where x1,y1 is your point